OpenStudy (anonymous):

Which expression is the simplified form of the square root of z^7 ?

OpenStudy (anonymous):

(z^7)^(1/2)= z^(7/2)

OpenStudy (anonymous):

it has to be one of these? z square root of z^3 z^2 square root of z z^3 square root of z z^3 square root of z^2

OpenStudy (phi):

Here's a hint \[z^{7/2}=z^{\frac{6}{2}+\frac{1}{2}}= z^{\frac{6}{2}}z^{\frac{1}{2}}\]

OpenStudy (anonymous):

is it the second one cause thats what i got the first time i did it.

OpenStudy (phi):

Look at the exponents in my hint. 6/2 =3 and 1/2 means square root.

OpenStudy (anonymous):

okay i am still a little cunfused but i think its the last one?

OpenStudy (anonymous):

try again

OpenStudy (phi):

\[z^{\frac{6}{2}}z^{\frac{1}{2}}= z^{3}z^{1/2}\] by definition, \[z^{1/2}= \sqrt{z}\]

OpenStudy (anonymous):

okays so the third one?

OpenStudy (phi):

Yes. Can you figure out why you are confused?

OpenStudy (anonymous):

yeah sorry this question just confused me more than the others thanks your your help! :)

OpenStudy (phi):

practice makes perfect.